If a garden has an area of 252 ft^2. Its length is 4 ft more than its width. What are the dimensions of the garden?
Answer by stanbon(75887)

If a garden has an area of 252 ft^2. Its length is 4 ft more than its width. What are the dimensions of the garden?
-----
width = w
length = w+4
------
Equation:
w(w+4) = 252
w^2 + 4w - 252 = 0
--------
width = 14 ft
length = w+4 = 18 ft
-------------
cheers,
Stan H.
---------